OpenStudy (anonymous):

solve for x: 3/7x + 2/9 = -4/7x - 1/5

OpenStudy (anonymous):

x = -19/45

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

You put all the xs on one side and the numbers on the other side of the equal sign. 3/7x+2/9+1/5=-4/7x 2/9+1/5=-4/7x-3/7x 2/9+1/5=-x 10/45+9/45=-x 19/45=-x x=-19/45

OpenStudy (anonymous):

Thanks!